OpenStudy (anonymous):

State what the length of the other leg is in a right triangle whose hypotenuse is 13 and one leg is 12 by using the Pythagorean Theorem. A. 4 B. 5 C. 3 D. 6

OpenStudy (adamk):

Do you know Pythagorean Theorem? It's a^2 + b^2 = c^2

OpenStudy (adamk):

So what number squared and added to 144 (12 squared) will give you 13^2?

OpenStudy (anonymous):

6? @AdamK

OpenStudy (adamk):

No, try again

OpenStudy (misty1212):

no dear you got a calculator? compute \[\sqrt{13^2-12^2}\] and see what you get

OpenStudy (adamk):

144 + 36 does not equal 169

OpenStudy (anonymous):

im gonna say 4 but i feel like im wrong

OpenStudy (adamk):

Please try to do the question. Do what @misty1212 said and you'll have the answer!

OpenStudy (anonymous):

5